OpenStudy (anonymous):

Simplify. Assume x represents a positive real number. The square root of 80x^4

OpenStudy (psymon):

\[\sqrt{80x ^{4}} = \sqrt{80}*\sqrt{x ^{4}}\] The exponent part isn't that bad, it's just whether or not you know how to simplify radical expressions.

OpenStudy (psymon):

For the exponent part, know that square root x is the same as \[x ^{\frac{ 1 }{ 2 }} \] So that means you have: \[\sqrt{80}*(x ^{4})^{\frac{ 1 }{ 2 }}\] Hopefully that makes simplifying the exponent part of it easier.